# Towards a general-purpose foundation model for fMRI analysis

### v1

@@
Cheng Wang1, Yu Jiang1, Zhihao Peng1, Chenxin Li1, Changbae Bang2, Lin Zhao3, Jinglei Lv4,
Jorge Sepulcre5, Carl Yang6, Lifang He7, Tianming Liu3, Daniel Barron8, Quanzheng Li9,
Randy Hirschtick9, Byung-Hoon Kim2, Xiang Li9,10∗, Yixuan Yuan1∗
@@
Massachusetts General Hospital, Boston, MA
@@ 1 Abstract
Functional Magnetic Resonance Imaging (fMRI) is crucial for studying brain function and diagnosing neurological disorders. However, existing methods for fMRI analysis suffer from reproducibility and transferability challenges, due to complex pre-processing pipelines and task-specific model designs. In this work, we introduce a Neuroimaging Foundation Model with Spatial-Temporal Optimized and Rpresentation Modeling (NeuroSTORM) that learns generalizable representations directly from 4D fMRI volumes, and achieves efficient pre-trained knowledge transfer across diverse downstream applications. Specifically, NeuroSTORM is pre-trained on a remarkable 28.65 million fMRI frames ($>$9,000 hours) collected from over 50,000 subjects, spanning multiple centers and covering ages 5 to 100. NeuroSTORM employs a shifted scanning strategy based on a Mamba backbone, enabling efficient direct processing of 4D fMRI volumes. Moreover, we introduce a spatial-temporal optimized pre-training strategy coupled with a task-specific prompt tuning technique, learning transferable fMRI features for efficient downstream adaptation. Experimental results show that NeuroSTORM consistently outperforms existing methods across five diverse downstream tasks: age and gender prediction, phenotype prediction, disease diagnosis, fMRI(-to-image) retrieval, and task-based fMRI (tfMRI) state classification. To further validate the clinical utility of NeuroSTORM, we evaluate it on two clinical datasets comprising patients with 17 different diagnoses from hospitals in the United States, South Korea, and Australia. NeuroSTORM maintains high relevance in predicting psychological/cognitive phenotypes and achieves the best disease diagnosis performance among all existing methods. In summary, NeuroSTORM offers a standardized, open-source foundation model to enhance reproducibility and transferability in fMRI analysis for clinical applications.
@@ 2 Introduction
Functional Magnetic Resonance Imaging (fMRI) is a non-invasive imaging technique widely used to study brain function and structure [1], and is also utilized in clinical practice for pre-surgical mapping, the assessment and monitoring of various neurological disorders [2, 3]. Moreover, it provides robust technological support for brain signal decoding and human-computer interaction in brain-computer interface research [4]. Driven by the advances in deep learning, recent works in fMRI analysis have achieved substantial improvements in the accuracy and efficiency of interpreting complex neural patterns from both resting-state fMRI (rsfMRI) and task fMRI (tfMRI) data  [5, 6, 7, 8, 9, 10, 11, 12]. Despite its theoretical and methodological progress, the field of fMRI remains fragmented across data formats, preprocessing pipelines, and analytic models, challenges recently summarized by [13]. Specifically, analytic reproducibility is fundamental to neuroimaging research, and overlooking it can introduce bias into subsequent analyses [14]. While prior studies related to reproducibility and transferability have focused on identifying reliable biomarkers [15], standardizing pipelines [16], and validating test-retest reliability [17], most of them remain tailored to specific fMRI applications and lack the generalizability to other tasks. Thus, there is an opportunity to develop a model that is intrinsically generalizable across diverse experimental settings and conditions, which can then enable both reproducibility and transferability within a unified framework.
@@
Figure 1: Overview of our proposed NeuroSTORM framework. (a) Data corpus and pre-processing: The model is pre-trained on a collection of publicly available datasets, including over 50,000 rsfMRI and 16,000 tfMRI sequences. All data are aligned to 2mm MNI152 space to create standardized 4D volumes. To facilitate the visualization of datasets with varying scales in the diagram, we use logarithmic values instead of raw numbers. (b) NeuroSTORM pre-training: The model utilizes a masked autoencoder paradigm with a STRD module to enhance the learning of long-range spatiotemporal relationships. (c) Downstream tasks with fine-tuning: For evaluation, NeuroSTORM employs a SWM Backbone and TPT techique for efficient fine-tuning to various downstream tasks. The benchmark includes age and gender prediction, phenotype prediction, disease diagnosis, fMRI(-to-image) retrieval, and tfMRI state classification. (d) Comprehensive performance evaluation: We systematically benchmark NeuroSTORM against previous state-of-the-art models across a diverse set of downstream tasks. The radar chart illustrates NeuroSTORM’s consistent performance improvements.
@@
In line with recent calls for scalable, generalizable frameworks for fMRI [13], we introduce the Neuroimaging Foundation Model with Spatial-Temporal Optimized and Rpresentation Modeling (NeuroSTORM), a general-purpose fMRI foundation model designed to enhance reproducibility and transferability through large-scale pre-training and architectural innovation (Fig. 1(a)). To enhance computational efficiency in 4D fMRI processing, we introduce a Shifted-Window Mamba (SWM) backbone, which combines linear-time state-space modeling with shifted-window mechanisms to reduce complexity and GPU memory usage. During pre-training, we propose a Spatiotemporal Redundancy Dropout (STRD) module (Fig. 1(b)) for effective learning of inherent characteristics in fMRI data, thereby improving the model’s robustness and reproducibility. For downstream task adaptation, our Task-specific Prompt Tuning (TPT) strategy (Fig. 1(c)) employs a minimal number of trainable, task-specific parameters when fine-tuning NeuroSTORM for new tasks. This provides a simple and integrated approach to applying NeuroSTORM across diverse applications. The corpus used to pre-train NeuroSTORM integrates four large-scale neuroimaging datasets: UK Biobank [26] (40,842 participants), ABCD [27] (9,448 children), and the HCP datasets [28] (HCP-YA, HCP-A, and HCP-D; totaling over 2,500 subjects). This represents the largest multi-source fMRI training dataset assembled to date. Spanning diverse demographics (ages 9–80), clinical conditions, and acquisition protocols, the corpus ensures broad biological and technical variation.
@@
To validate the performance and transferability of NeuroSTORM, we established a comprehensive fMRI analysis benchmark, including five downstream tasks: age and (reported) gender prediction, phenotype prediction, disease diagnosis, fMRI retrieval, and task-based fMRI (tfMRI) state classification. Notably, we assessed the clinical applicability of NeuroSTORM on two clinical datasets from hospitals in both the United States and Australia, Transdiagnostic Connectome Project (TCP) [29] and Motor Neuron Disease (MND) [30]. The TCP dataset consists of 245 participants from Brain Imaging Center of Yale University or McLean Hospital in the United States, including both healthy controls and individuals covering a range of psychiatric disorders. The MND dataset includes 36 participants diagnosed with Amyotrophic Lateral Sclerosis and 23 controls, collected at the Royal Brisbane and Women’s Hospital in Australia. NeuroSTORM outperforms or matches state-of-the-art models across all five tasks, demonstrating strong transferability in diverse applications. Moreover, we simulated a data-scarce scenario by limiting the proportion of fine-tuning data in order to evaluate data utilization efficiency. NeuroSTORM demonstrates minor performance degradation on most datasets. Notably, we evaluated NeuroSTORM’s ability to predict clinical phenotypes and perform disease classification in clinical datasets, with experimental results highlighting its clinical value and transferability.
@@ 3 Results
The downstream task of age and (reported) gender prediction utilizes HCP-YA [28], HCP-A, HCP-D, UKB [26] and ABCD [27] datasets. Phenotype prediction is based on HCP-YA and TCP [29] dataset. Disease diagnosis draws on multiple datasets, including HCP-EP [28], ABIDE [31], ADHD200 [32], COBRE [33], UCLA [34] and MND [30]. The fMRI retrieval task is evaluated on the NSD [35] dataset. Task-based fMRI (tfMRI) state classification task uses the HCP-YA dataset. No annotations are included in the NeuroSTORM pre-training corpus. In all experiments, the downstream task datasets are split into training, validation, and testing sets in an 8:1:1 ratio. Additionally, we evaluate NeuroSTORM’s robustness in data-scarce scenarios by employing varying percentages of fine-tuning data in the datasets. For fMRI retrieval, we conduct experiments using the LAION-5B [36] dataset as the retrieval candidate pool.
@@ 3.1 Age and Gender Prediction
Figure 2: Evaluation of NeuroSTORM’s performance in (reported) gender classification and age regression. (a) Gender classification performance: NeuroSTORM consistently outperformed competing ROI-based and volume-based methods across datasets, including HCP-YA, HCP-A, HCP-D, UKB and ABCD. It achieved the highest accuracy and area under the curve (AUC) metrics. (b) Age regression error: In age regression tasks, NeuroSTORM achieved the lowest error rates, surpassing competing methods across multiple datasets. (c) Performance in data-scarce scenarios: NeuroSTORM demonstrated strong adaptability when trained with limited data. Remarkably, even with only 10%-50% of the training data, it maintained competitive performance in both age and gender prediction tasks. With increasing training data proportions, its performance steadily improved, achieving optimal results with full datasets. The variance in the result images was computed based on three independent experimental replicates to ensure statistical reliability.
To evaluate the reproducibility and transferability of NeuroSTORM, we first assessed its performance in predicting age and (reported) gender from rsfMRI sequences compared to state-of-the-art fMRI analysis methods. Age and gender are critical sociodemographic variables correlating with structural brain changes, making this a fundamental fMRI analysis task. NeuroSTORM consistently outperformed both ROI-based methods (BrainGNN [37], BNT [38]) and volume-based approaches across all benchmark datasets. On the HCP-YA [28] dataset, our model achieved gender classification accuracy of 93.3% (AUC 97.6%) versus Com-brainTF’s 85.6% accuracy (AUC 94.7%). Similar superiority was observed in HCP-A (92.3% vs 85.1% accuracy) and HCP-D (85.4% vs 82.6% accuracy). For age prediction on ABCD, NeuroSTORM reduced Mean Square Error (MSE) to 6.9 compared to SwiFT [24]’s 7.6. While traditional ROI-based methods showed advantages in small datasets due to lower dimensionality and higher signal-to-noise ratios, NeuroSTORM’s architecture ultimately surpassed them through effective integration of structural and functional features (Fig. 2).
NeuroSTORM demonstrated exceptional efficiency in data-scarce scenarios. With only 50% fine-tuning data on HCP-D, it maintained 83.3% accuracy compared to TFF’s 82.5%. The model’s pre-training strategy enabled effective knowledge transfer to UKB dataset predictions, achieving 99.2% accuracy with just 30% fine-tuning data versus SwiFT’s 98.9%. This transferability advantage persisted across multiple domains, with NeuroSTORM maintaining less than 5% performance degradation when trained on 30% of ABCD data compared to more than 12% degradation observed in volume-based baselines.
@@ 3.2 Phenotype Prediction
Figure 3: Performance evaluation of NeuroSTORM in phenotype prediction tasks. (a) HCP-YA dataset: NeuroSTORM demonstrates superior Pearson Correlation Coefficients (PCC) across diverse phenotype scores, including MMSE Score (P01), Social Task Performance (P02), Cognitive Total Score (Age Adjusted) (P03), Emotion Task Accuracy (P04), Language Task Accuracy (P05), and Strength Score (Age Adjusted) (P06), significantly outperforming ROI-based and volume-based methods including BrainGNN, BNT, and SwiFT. Even in data-scarce scenarios (10%-50%), NeuroSTORM maintains competitive PCC performance. The variance in the result images was computed based on three independent experimental replicates to ensure statistical reliability. (b) TCP dataset: NeuroSTORM is evaluated for its ability to predict crucial disease-related scores in subjects with psychiatric disorders, including Anxiety Sensitivity (P07), CGI Severity Score (P08), DASS Anxiety Score (P09), DASS Stress Score (P10), PANSS General Score (P11), PANSS Negative Symptoms (P12), PANSS Positive Symptoms (P13), NEO Agreeableness Score (P14), TCI Harm Avoidance Score (P15), and TCI Cautiousness Score (P16). The model achieves positive correlations across multiple phenotype scores.
@@
As shown in Fig. 3, NeuroSTORM demonstrated improved performance compared to ROI-based and volume-based methods across various scores. For Cognitive Total Score (P03), it achieved a Pearson Correlation Coefficient (PCC) of 0.378 in the MMSE Score (P01), compared to 0.301 (ROI-based) and 0.347 (volume-based baselines). For Emotion Task Accuracy (P04), NeuroSTORM attained a PCC of 0.429, which was 6.8–12.3% higher than alternative approaches. Further analysis indicated consistent performance in capturing functional patterns. To assess data efficiency, training data was progressively reduced from 100% to 10%. With 10% labeled data, the model retained 73.5% of its full-data MMSE (P01) performance (PCC=0.278 vs 0.378) and 69.9% in Emotion Task (P04) (PCC=0.300 vs 0.429). Performance increased with data availability, reaching 91.5% (MMSE PCC=0.348) and 88.3% (Emotion Task PCC=0.379) efficacy at 50% training data. This improvement trend was observed across all scores, including Strength Score (P06) (from PCC=0.208 at 10% to 0.297 at 100%) and Social Task (P02) (0.189 to 0.249). The Mean Absolute Error (MAE) metrics with limited data showed similar scaling trends, indicating the model’s effectiveness in data-scarce scenarios.
For the TCP dataset, NeuroSTORM was evaluated on its ability to predict clinical phenotypes associated with psychiatric disorders. The model’s performance varied across different phenotypes, achieving PCC ranging from 0.202 to 0.587. Specifically, higher correlations were observed for PANSS Positive Symptoms (P13) (PCC=0.559) and PANSS Negative Symptoms (P12) (PCC=0.464) scores, while lower correlations were noted for Anxiety Sensitivity (P07) (PCC=0.234) and TCI Cautiousness Score (P16) (PCC=0.295). The MAE across tasks varied from 1.455 to 13.03, reflecting differences in prediction accuracy among the phenotypes. These results illustrate NeuroSTORM’s capacity to predict a range of clinical phenotypes, with performance differing based on the specific measures being assessed.

### v2

@@
Cheng Wang1, Yu Jiang1, Zhihao Peng1, Chenxin Li1, Changbae Bang2, Lin Zhao3, Wanyi Fu4,
Jinglei Lv5, Jorge Sepulcre6, Carl Yang7, Lifang He8, Tianming Liu3, Xue-Jun Kong9,
Quanzheng Li10, Daniel S. Barron11, Anqi Qiu12, Randy Hirschtick10, Byung-Hoon Kim2,
Hongbin Han4∗,Xiang Li10,13∗, Yixuan Yuan1∗
@@
Peking University Health Science Center, Beijing, China
@@
Boston Children’s Hospital, Boston, MA
Massachusetts General Hospital, Boston, MA
@@
Hong Kong Polytechnic University, Hong Kong
@@
11footnotetext: Cheng Wang; Yu Jiang; Zhihao Peng; Chenxin Li; Changbae Bang; Lin Zhao; Wanyi Fu; Jinglei Lv; Jorge Sepulcre; Carl Yang; Lifang He; Tianming Liu; Xue-Jun Kong; Quanzheng Li; Daniel Barron; Anqi Qiu; Randy Hirschtick; Byung-Hoon Kim; Hongbin Han: hanhongbin@bjmu.edu.cn; Xiang Li: xli60@mgh.harvard.edu; Yixuan Yuan: yxyuan@$e$.cuhk.edu.hk.
@@ 1 Abstract
Functional MRI (fMRI) is crucial for studying brain function and diagnosing neurological disorders. However, existing analysis methods suffer from reproducibility and transferability challenges due to complex preprocessing pipelines and task-specific model designs. In this work, we introduce a Neuroimaging Foundation Model with Spatial-Temporal Optimized and Representation Modeling (NeuroSTORM) that learns generalizable representations directly from 4D fMRI volumes and enables efficient transfer to diverse downstream applications. Specifically, NeuroSTORM is pre-trained on 28.65 million fMRI frames from over 50,000 subjects, spanning multiple centers and ages 5 to 100. It combines an efficient spatiotemporal modeling design and lightweight task adaptation to enable scalable pre-training and fast transfer to downstream applications. Here we show that NeuroSTORM consistently outperforms existing methods across five downstream tasks, including demographic prediction, phenotype prediction, disease diagnosis, re-identification, and state classification. On two multi-hospital clinical cohorts with 17 diagnoses, NeuroSTORM achieves the best diagnosis performance while remaining predictive of psychological and cognitive phenotypes. These results suggest that NeuroSTORM could become a standardized foundation model for reproducible and transferable fMRI analysis.
@@ 2 Introduction
Functional Magnetic Resonance Imaging (fMRI) is a non-invasive imaging technique widely used to study brain function and structure [1], and is also investigated in research contexts for potential use in assessment and monitoring of various neurological disorders [2, 3]. Moreover, it provides robust technological support for brain signal decoding and human-computer interaction in brain-computer interface research [4]. Driven by the advances in deep learning, recent works in fMRI analysis have achieved substantial improvements in the accuracy and efficiency of interpreting complex neural patterns from both resting-state fMRI (rsfMRI) and task fMRI (tfMRI) data [5, 6, 7, 8, 9, 10, 11, 12]. Despite its theoretical and methodological progress, the field of fMRI remains fragmented across data formats, preprocessing pipelines, and analytic models, challenges recently summarized by [13]. Specifically, analytic reproducibility is fundamental to neuroimaging research, and overlooking it can introduce bias into subsequent analyses [14]. While prior studies related to reproducibility and transferability have focused on identifying reliable biomarkers [15], standardizing pipelines [16], and validating test-retest reliability [17], most of them remain tailored to specific fMRI applications and lack the generalizability to other tasks. Thus, there is an opportunity to develop a model that is intrinsically generalizable across diverse experimental settings and conditions, which can then enable both reproducibility and transferability within a unified framework.
Figure 1: Overview of our proposed NeuroSTORM framework. (a) Data corpus and preprocessing: The model is pre-trained on a collection of publicly available datasets, including over 50,000 rsfMRI and 16,000 tfMRI sequences. All data are aligned to 2mm MNI152 space to create standardized 4D volumes. Since we did not collect the complete metadata for the DMT-HAR-MED dataset, DMT-HAR-MED is not included in statistical information of the datasets. (b) NeuroSTORM pre-training: The model utilizes a masked autoencoder paradigm with a STRD module to enhance the learning of long-range spatiotemporal relationships. (c) Downstream tasks with fine-tuning: For evaluation, NeuroSTORM employs a SWM Backbone and TPT techique for efficient fine-tuning to various downstream tasks. The benchmark includes age and gender prediction, phenotype prediction, disease diagnosis, fMRI re-identification, and fMRI state classification. (d) Comprehensive performance evaluation: We systematically benchmark NeuroSTORM against previous ROI-based and Volume-based state-of-the-art models across a diverse set of downstream tasks. The radar chart illustrates NeuroSTORM’s consistent performance improvements.
@@
In line with recent calls for scalable, generalizable frameworks for fMRI [13], we introduce the Neuroimaging Foundation Model with Spatial-Temporal Optimized and Rpresentation Modeling (NeuroSTORM), a general-purpose fMRI foundation model designed to enhance reproducibility and transferability through large-scale pre-training and architectural innovation (Fig. 1(a)). To enhance computational efficiency in 4D fMRI processing, we introduce a Shifted-Window Mamba (SWM) backbone, which combines linear-time state-space modeling with shifted-window mechanisms to reduce complexity and GPU memory usage. During pre-training, we propose a Spatiotemporal Redundancy Dropout (STRD) module (Fig. 1(b)) for effective learning of inherent characteristics in fMRI data, thereby improving the model’s robustness and reproducibility. For downstream task adaptation, our Task-specific Prompt Tuning (TPT) strategy (Fig. 1(c)) employs a minimal number of trainable, task-specific parameters when fine-tuning NeuroSTORM for new tasks. This provides a simple and integrated approach to applying NeuroSTORM across diverse applications. The corpus used to pre-train NeuroSTORM integrates three large-scale neuroimaging datasets: UK Biobank [26] (40,842 participants), ABCD [27] (9,449 children), and the HCP datasets [28] (HCP-YA, HCP-A, and HCP-D; totaling over 2,300 subjects). Spanning diverse demographics (ages 9-100), clinical conditions, and acquisition protocols, the corpus ensures broad biological and technical variation.
@@
To validate the performance and transferability of NeuroSTORM, we established a comprehensive fMRI analysis benchmark, including five downstream tasks: age and (reported) gender prediction, phenotype prediction, disease diagnosis, fMRI re-identification, and fMRI state classification. Notably, we assessed the clinical applicability of NeuroSTORM on two clinical datasets from hospitals in both the United States and Australia, Transdiagnostic Connectome Project (TCP) [29] and Motor Neuron Disease (MND) [30]. The TCP dataset consists of 245 participants from Brain Imaging Center of Yale University or McLean Hospital in the United States, including both healthy controls and individuals covering a range of psychiatric disorders. The MND dataset includes 36 participants diagnosed with Amyotrophic Lateral Sclerosis and 23 controls, collected at the Royal Brisbane and Women’s Hospital in Australia. NeuroSTORM outperforms or matches state-of-the-art models across all five tasks, demonstrating strong transferability in diverse applications. Moreover, we simulate a data-scarce scenario by limiting the proportion of fine-tuning data in order to evaluate label efficiency. NeuroSTORM demonstrates minor performance degradation on most datasets. Notably, we evaluated NeuroSTORM’s ability to predict clinical phenotypes and perform disease classification in clinical datasets, with experimental results highlighting its clinical value and transferability.
@@ 3 Results
The downstream task of age and (reported) gender prediction utilizes HCP-YA [28], HCP-A, HCP-D, UKB [26] and ABCD [27] datasets. Phenotype prediction is based on HCP-YA and TCP [29] dataset. Disease diagnosis draws on multiple datasets, including HCP-EP [28], ADHD200 [31], COBRE [32], UCLA [33] and MND [30]. The fMRI re-identification task is evaluated on the HCP-YA. fMRI state classification task uses the HCP-YA and DMT-HAR-MED [34] dataset. No annotations are included in the NeuroSTORM pre-training corpus. We compare NeuroSTORM against with baselines from two major paradigms: (i) ROI-based methods that convert the 4D fMRI into 2D region-of-interest (ROI) connectomes, includes BrainGNN [35], BrainNet Transformer (BNT) [36], LG-GNN [37], Com-brainTF [38], IBGNN [39], Brain-JEPA [40] and BrainLM [8]; and (ii) volume-based methods that directly precess voxel-wise 4D volumes, TFF [25] and SwiFT [24]. SwiFT is particularly relevant to our work as it also employs an end-to-end backbone tailored for spatio-temporal 4D inputs. All baselines are re-implemented or run with authors’ code under identical preprocessing to ensure a fair comparison. In all experiments, the downstream task datasets are split into training, validation, and testing sets in an 8:1:1 ratio. The same dataset split and training details are applied to all comparison methods. Additionally, we evaluate NeuroSTORM’s label efficiency in data-scarce scenarios by employing varying percentages of fine-tuning data in the datasets.
Figure 2: Evaluation of NeuroSTORM’s performance in (reported) gender classification and age regression tasks. (a) Gender classification performance: NeuroSTORM consistently outperformed competing ROI-based and volume-based methods across datasets, including HCP-YA, HCP-A, HCP-D, UKB and ABCD. (b) Age regression error: In age regression tasks, NeuroSTORM achieved the lowest error rates, surpassing competing methods across multiple datasets. (c) Label efficiency results: NeuroSTORM demonstrated strong adaptability when trained with limited data. Remarkably, even with only 10%-50% of the training data, it maintained competitive performance in both age and gender prediction tasks. With increasing training data proportions, its performance steadily improved, achieving optimal results with full datasets. Variances were estimated from five technical replicates.Pairwise significance markers were computed using a two-sided paired t-test (ttest_rel) without multiple-comparison correction, and the corresponding P-value is annotated.
@@ 3.1 Age and Gender Prediction
To evaluate the transferability of NeuroSTORM, we first assessed its performance in predicting age and gender from rsfMRI sequences compared to state-of-the-art fMRI analysis methods. Age and gender are fundamental sociodemographic variables correlating with structural brain changes, making this a fundamental fMRI analysis task. Since the dataset metadata includes only reported gender, the prediction target in this downstream task is that label. NeuroSTORM consistently outperformed both ROI-based methods and volume-based approaches across all benchmark datasets. On HCP-YA [28] it reached a gender-classification accuracy of 93.28%, surpassing the best volume-based method SwiFT [24] at 92.66% and the best ROI-based method Brain-JEPA at 91.80%. Similar superiority is observed in HCP-A (92.16% vs 91.78% for SwiFT and 89.24% for Brain-JEPA) and HCP-D (85.38% vs 84.18% for SwiFT and 83.04% for Brain-JEPA). For age prediction on ABCD, NeuroSTORM reduced Mean Square Error (MSE) to 7.0 compared to SwiFT [24]’s 7.72 (best volume baseline) and LG-GNN’s 10.40 (best ROI baseline). While traditional ROI-based methods showed advantages in small datasets due to lower dimensionality and higher signal-to-noise ratios, NeuroSTORM’s architecture ultimately surpassed them through effective integration of structural and functional features (Fig. 2).
NeuroSTORM also exhibited strong label efficiency in data-scarce scenarios. For instance, on HCP-D, training with only 50% of the available labels still produced 83.7% accuracy, clearly ahead of SwiFT’s 80.1% and BrainLM’s 77.0%. When the training set was limited to 30%, NeuroSTORM retained 83.2% accuracy, losing less than 3% absolute performance relative to full-data training, whereas SwiFT and BrainLM dropped by more than 5% and 6%, respectively. These results show that the pre-training strategy of NeuroSTORM enables efficient knowledge transfer and robust generalization even when labeled data are scarce.
@@ 3.2 Phenotype Prediction
Figure 3: Performance evaluation of NeuroSTORM in phenotype prediction task. (a) HCP-YA dataset: NeuroSTORM demonstrates superior Pearson Correlation Coefficients (PCC) across diverse phenotype scores, including MMSE Score (P01), Social Task Performance (P02), Cognitive Total Score (Age Adjusted) (P03), Emotion Task Accuracy (P04), Language Task Accuracy (P05), and Strength Score (Age Adjusted) (P06). Even in data-scarce scenarios (10%-50%), NeuroSTORM maintains competitive PCC performance. (b) TCP dataset: NeuroSTORM is evaluated for its ability to predict crucial disease-related scores in subjects with psychiatric disorders, including Anxiety Sensitivity (P07), CGI Severity Score (P08), DASS Anxiety Score (P09), DASS Stress Score (P10), PANSS General Score (P11), PANSS Negative Symptoms (P12), PANSS Positive Symptoms (P13), NEO Agreeableness Score (P14), TCI Harm Avoidance Score (P15), and TCI Cautiousness Score (P16). Variances were estimated from five technical replicates. Pairwise significance markers were computed using a two-sided paired t-test (ttest_rel) without multiple-comparison correction, and the corresponding P-value is annotated.
As shown in Fig. 3, NeuroSTORM demonstrated improved performance compared to ROI-based and volume-based methods across various scores. For MMSE (P01) it reaches 0.372, which is 9.1% higher than the strongest volume baseline SwiFT (0.341) and 31.5% higher than the best ROI baseline Brain-JEPA (0.336). Consistent gains are observed for Social Task Performance (P02, 0.247 vs 0.227 for SwiFT), Cognitive Total Score (P03, 0.328 vs 0.317 for SwiFT), Emotion Task Accuracy (P04, 0.419 vs 0.373 for SwiFT), Language Task Accuracy (P05, 0.284 vs 0.254 for SwiFT) and Strength Score (P06, 0.297 vs 0.260 for TFF). Label-efficiency curves reveal a slower degradation for NeuroSTORM: on P01 it loses only 7.8% when the label budget drops from 100% to 50%, whereas SwiFT and BrainLM drop 13.5% and 33.1%, respectively. NeuroSTORM reaches 0.348 PCC, surpassing SwiFT trained with the full dataset (0.341). For P04, NeuroSTORM attains 0.375 with 30% labels, which is close to SwiFT’s 0.373 with 100%, and rises to 0.381 with 50% labels, illustrating a consistently flatter performance curve than competing methods.
For the TCP dataset, NeuroSTORM was evaluated on its ability to predict clinical phenotypes associated with psychiatric disorders. The model’s performance varied across different phenotypes, achieving PCC ranging from 0.234 to 0.558. Specifically, highest correlations are obtained for PANSS Positive Symptoms (P13, PCC=0.558) and PANSS Negative Symptoms (P12, PCC=0.464). Lower correlations are observed for Anxiety Sensitivity (P07, PCC=0.234) and TCI Cautiousness (P16, PCC=0.296). Mean-absolute error ranges from 1.455 (P13) to 13.23 (P15), reflecting the varying difficulty of the prediction tasks. These results confirm that NeuroSTORM can model a broad spectrum of clinically relevant traits, though performance naturally varies with the specific target.
Figure 4: Evaluation of NeuroSTORM’s performance on disease diagnosis task. (a) Classification accuracy on ADHD200, COBRE, HCP-EP, MND and UCLA shows that NeuroSTORM consistently outperforms all ROI-based and volume-based baselines, highlighting its strong generalisability across neurological and psychiatric disorders. (b) Label efficiency analysis demonstrates that NeuroSTORM maintains robust performance even when trained with only a fraction of the fine-tuning data, underscoring its suitability for data-scarce scenarios. Variances were estimated from five technical replicates. Pairwise significance markers were computed using a two-sided paired t-test (ttest_rel) without multiple-comparison correction, and the corresponding P-value is annotated.